What are the slope and y-intercepts of x=-9?
Answer by Nate(3500)

x = -9 ...
Simplify the meaning by picking multiple coordinates.
( x , -2 ), ( x , -1 ), ( x , 0 ), ( x , 1 ), ( x , 2 )
There is y to plug in these values: x = 0y - 9
( -9 , -2 ), ( -9 , -1 ), ( -9 , 0 ), ( -9 , 1 ), ( -9 , 2 )
This is a vertical line, so the line doesn't intersect with the y-axis ... no y-intercepts.
slope = (y2 - y1) / (x2 - x1) = (2 - 1) / (-9 + 9) = 1 / 0
The slope is undefined.